In a hydrogen atom, an electron with a charge of (-)1.6E-19 orbits a proton with the same but positive charge at a distance 65 picometers away (that's trillionths of a meter or 10-12 m). What electric force F, in N, binds the two particles together? Use "E" notation for your answer.
Answer #1

Solution:

Using the coulomb's law,

F = kQ2/r2

= 8.99*109 *[1.60*10-19]2/[65*10-12]2

= 5.45*10-8 N

= 5.45 E-18 N.

I hope you understood the problem and got your answers, If yes rate me!! or else comment for a better solutions